# Time Evidence Fusion Network: Multi-source View in Long-Term Time Series Forecasting

### v1

@@ Abstract
In real-world scenarios, time series forecasting often demands timeliness, making research on model backbones a perennially hot topic. To meet these performance demands, we propose a novel backbone from the perspective of information fusion. Introducing the Basic Probability Assignment (BPA) Module and the Time Evidence Fusion Network (TEFN), based on evidence theory, allows us to achieve superior performance. On the other hand, the perspective of multi-source information fusion effectively improves the accuracy of forecasting. Due to the fact that BPA is generated by fuzzy theory, TEFN also has considerable interpretability. In real data experiments, the TEFN partially achieved state-of-the-art, with low errors comparable to PatchTST, and operating efficiency surpass performance models such as Dlinear. Meanwhile, TEFN has high robustness and small error fluctuations in the random hyperparameter selection. TEFN is not a model that achieves the ultimate in single aspect, but a model that balances performance, accuracy, stability, and interpretability. Code is available at this repository: https://github.com/ztxtech/Time-Evidence-Fusion-Network.
1 Introduction
The changes of things are accompanied by the passage of time, and time series appear in more and more application scenarios, such as health diagnosis, financial analysis, weather forecasting, and so on. Optimizing structure of backbone is an effective way to improve the performance of time series forecasting, where linear-based models such as Dlinear often outperform transformer-based models such as Crossformer [18, 20]. Therefore, we hope to propose a model that balances performance, prediction accuracy, stability, and interpretability to provide support for highly reliable scenarios with time requirements.
Evidence theory is a model of uncertain reasoning commonly used for decision-making in the fusion of multiple information [14, 6]. Meanwhile, evidence theory is often appeared in reasoning framework based on symbolic logic, with high interpretability. Denoeux has implemented evidence neural networks for classification tasks [8]. Therefore, we propose the Time Evidence Fusion Network (TEFN) for regression tasks. The strategy adopted by TEFN is to divide the time series into two perspectives: time and channels, treating each channel/time as an independent information source, and using an information fusion framework for prediction in Figure 1. Therefore, TEFN is a simple and naive model, without any technical modifications, and is a new backbone.
@@ 1 Introduction
Figure 1: The channel dimension and time dimension of multivariate time series can be considered as different sources of information. Firstly, merge information sources of the same type, and finally integrate information sources from both channel and time dimensions to complete the final prediction.
In TEFN, we propose a novel module Basic Probability Assignment (BPA) Module. Based on the evidence theory of decision-making process, BPA maps different sources of information to distributions related to the target outcome. The target distribution is achieved by the fuzzy membership of different fuzzy sets [7]. In image processing, convolution condenses the required features of the target [4]. Unlike convolution, BPA is an expansion process in Figure 2. On the contrary, BPA considers all possibilities in the event space composed of subsets of the sample space, expanding on different events, which is more effective for simple structures like time series and can fully explore the different hidden information inside. For different distributions generated by multiple information sources, we consider sources as samples and use the expected forecasting values as the final result.
@@
Figure 2: Comparison between convolution and basic probability assignment: convolution condenses information while basic probability assignment diverges to consider different possibilities.
In this article, we focus on the practical application details of TEFN, including its accuracy, memory usage, training speed, parameter sensitivity, and interpretability. At the same time, in order to ensure fairness, the experimental data and configuration files for comparison are all from open-source projects. TEFN achieves state-of-the-art (SOTA) on multiple real datasets, while training time and model parameters are much smaller than models based on Transformer [20] or Patch [12]. For random hyperparameter selection, TEFN exhibits weak fluctuations and is a stable model. The contributions of this article are as follows:
@@
  • This article proposes a new evidential backbone for extracting potential distributions of simple data structures.
@@
  • A neural network TEFN based on BPA for time series forecasting has been proposed, which leads previous research in accuracy, efficiency, stability, and interpretability.
@@
  • This article explicitly introduces relevant concepts from evidence theory, such as mass function, into the structure of neural networks, providing a new implementation approach for uncertainty reasoning.
2 Related Work
2.1 Dempster-Shafer Theory
@@
Dempster-Shafer theory, also known as Evidence Theory, operates on a different framework compared to traditional probability theory [14, 6]. It relies on the concept of mass functions, which provide a more flexible representation of uncertainty, accommodating weaker constraints than traditional probabilities. When combining two mass functions, denoted as $m_{1}$ and $m_{2}$, the resulting fused mass function $m$ is determined using the Dempster-Shafer Rule (DSR), as described in Equation 1.
$m(A)=\sum_{B\cap C=A}m_{1}(B)*m_{2}(C)$ (1)
2.2 Evidence Decision Making
@@
The process of Evidence Decision Making can be delineated into several sequential steps: basic probability assignment (BPA), evidence fusion, and decision making (DM) [7]. BPA serves as a pivotal initial phase, constituting a mapping from raw data to mass distribution, akin to the path from raw data to probability distribution. Subsequently, evidence fusion integrates distinct mass distributions from various data channels, culminating in a rationalized mass distribution. Finally, decision making transpires as the progression from the attained distribution towards the designated target.
3 Time Evidence Fusion Network
The structure of TEFN is naive, and well performance can be achieved through a structure with only $2$ BPA module similar to only $2$ convolution kernels. We believe that the new architecture of TEFN is more like a pure backbone model. The overall structure of TEFN is in Figure 3.
Figure 3: The overall structure of TEFN
3.1 Time Normalization and De-normalization
In order to accelerate converging to local optima more quickly, we referred to the normalization methods of Stationary in TEFN [11]. The normalization process calculates the mean $\mu$ in Equation 2 and variance difference $\sigma^{2}$ in Equation 3 for each piece of the time series $x$. The normalization and de-normalization processes are a pair of symmetric operations in Equation 4 and Equation 5, where $Net(\cdot)$ is the neural network operation performed on the normalized time series $x_{Norm}$.
$\mu=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}$ (2)
$\sigma^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}(x_{i}-\mu)^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}^{2}-\mu^{2}$ (3)
$x_{Norm}=\frac{x-\mu}{\sigma}$ (4)
$x_{De-Norm}=\sigma*\hat{y}_{Norm}+\mu=\sigma*Net(x_{Norm})+\mu$ (5)
3.2 Time Dimension Projection
In TEFN, a linear layer is used to project the length $L_{in}$ of the time dimension to length $L_{in}+L_{pred}$ where $L_{pred}$ is prediction length of current target.
3.3 Basic Probability Assignment Module
@@
Figure 4: The operation process of a BPA module for a time series $T$ of length $L=5$. The sample space size in the BPA module is set as $|S|=2$, and the event space size is set as $|2^{S}|=2^{|S|}=4$. The operation of the BPA module is similar to that of the convolution module, but the BPA module diffuses different masses instead of summing them in the convolution.
According to Evidence Theory, each dimension of a time series can be reduced to a mass function in the same set of event spaces. The event space of evidence theory usually chooses power sets $2^{S}$ of sample space $S$. If it is a classification problem, the sample space is all the labels. For regression problems, the sample space needs to be treated as anonymous samples, where the samples are mutually exclusive. Similarly, whether it is a classification or regression problem, each dimension of the data will generate a mass distribution in the event space. Fuzzy numbers commonly used in the generation of mass distribution of evidence.
Here we use a parameterized fuzzy membership $\mu$ in Equation 6, where $D$ (time dimension $T$ and channel dimension $C$ are available) is the dimension that generates the mass function, $j$ is corresponding to dimension $D$ and $i$ is corresponding to the other one. Parameter $w$ and $b$ represent the slope and intercept of the membership function, and $k$ is corresponding to event dimension $F$.
$m_{D,i,j,k}=\mu(x_{Norm,i,j})=w_{D,j,k}*x_{Norm,i,j}+b_{D,j,k}$ (6)
3.4 Expectation Fusion
For multivariate time series in shape $|T|*|C|$, there are two dimensions: time dimension $T$ and channel dimension $C$. So, the different mass distributions $m_{C},m_{T}$ can be generated simultaneously. And the fusion method here uses expectations in Equation 7 where $y_{j,k}$ is parameter and $m^{\prime}_{i,j,k}$ is corresponding mass function. Summing part $y_{j,k}*m^{\prime}_{i,j,k}$ can be regarded as a linear transform of $m_{i,j,k}$, and expectation is calculated by summing $m_{i,j,k}$ in Equation 7. There are two reasons why DSR are not used:
  1. 1.
    The mathematical operation corresponding to DSR is the point-to-point multiplication of mass distributions within the same dimension, which has a high computational complexity. At the same time, if the dimension is too high, it will lead to the disappearance of accuracy within a finite precision.
  2. 2.
    DSR is sensitive to one extreme distribution, such as a distribution with only one element being $1$, which can result in the final fusion result of the entire mass being directly equal to that distribution.
$\hat{y}_{Norm,D,i,j}=E_{D,F}(y_{j,k})=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}y_{j,k}*m^{\prime}_{i,j,k}=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}m_{i,j,k}$ (7)
At the end, results of time dimension $T$ and channel dimension $C$ will aggregate in Equation 8 where $\hat{y}_{norm}$ is normalized predicted results. The final prediction results are restored through the de-normalization layer.
$\hat{y}_{norm}=\sum_{D\in\{T,C\}}\hat{y}_{Norm,D}=\hat{y}_{Norm,T}+\hat{y}_{Norm,C}$ (8)
4 Experiments
Dataset
We will use TEFN for long-term prediction tasks. Table 1 shows the basic settings used in the experiment, including the datasets. In terms of datasets, we have selected $5$ common time series datasets, including Electricity [1], ETT (4 subsets) [10], Exchange [17], Traffic [2], Weather [3]. More dataset features are in the Appendix Table 5.
Table 1: Datasets and Basic Settings
Tasks
Datasets
Metrics
Series Length
Long-term
ETT (4 subsets), Electricity,
MSE, MAE 96$\sim$720
Forecasting
Traffic, Weather, Exchange
Baseline
The comparative experimental models include PatchTST (ICLR 2023) [12], Crossformer (ICLR 2023) [20], TiDE [5], TimeNet (ICLR 2023) [16], ETSformer [15], LightTS [19], Dlinear (AAAI 2023) [18], FEDformer (ICML 2022) [21], Stationary (NeurIPS 2022) [11] and Autoformer (NeurIPS 2021) [17] . The configurations of the comparative experiments are from the Time Series Library (https://github.com/thuml/Time-Series-Library). More implementation details are in the Appendix Section A.
4.1 Forecasting Result
TEFN partially achieved SOTA performance in long-term forecasting target in Table 2. It is worth noting that in the case of a simple structure with tiny parameters in Section 4.4, achieved approximate or even better performance than previous SOTA models with larger parameters, such as PatchTST [12]. Full forecasting results are in the Appendix Table 7.
Table 2: Prediction error. All results are the average of four predicted lengths $\{96,192,336,720\}$.

### v2

@@ Abstract.
In practical scenarios, time series forecasting necessitates timeliness, especially when dealing with large datasets. Consequently, the exploration of model architectures remains a perennially trending topic in research. To meet these performance demands, we propose a novel backbone from the perspective of information fusion. Introducing the Basic Probability Assignment (BPA) Module and the Time Evidence Fusion Network (TEFN), based on evidence theory, allows us to achieve superior performance. On the other hand, the perspective of multi-source information fusion effectively improves the accuracy of forecasting. Due to the fact that BPA is generated by fuzzy theory, TEFN also has considerable interpretability. In real data experiments, the TEFN partially achieved state-of-the-art, with low errors comparable to PatchTST, and operating efficiency surpass performance models such as Dlinear. Meanwhile, TEFN has high robustness and small error fluctuations in the random hyperparameter selection. TEFN is not a model that achieves the ultimate in single aspect, but a model that balances performance, accuracy, stability, and interpretability.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/ztxtech/Time-Evidence-Fusion-Network.
1. Introduction
The evolution of various phenomena is inherently tied to the progression of time, leading to an increasing prevalence of time series data in a multitude of application domains. From health diagnostics, where the progression of diseases is tracked over time, to financial analysis, where market trends are monitored for investment strategies, and weather forecasting, where atmospheric changes are predicted to safeguard communities, the importance of time series forecasting cannot be overstated. However, as the volume of data grows exponentially, time series forecasting faces significant performance challenges. The sheer scale of big data necessitates more efficient and faster algorithms to process and predict within acceptable time frames, which has become a pressing issue in the field.
@@ 1. Introduction
Figure 1. The channel dimension and time dimension of multivariate time series can be considered as different sources of information. Firstly, merge information sources of the same type, and finally integrate information sources from both channel and time dimensions to complete the final prediction.
In response to these challenges, researchers have been focusing on optimizing the structure of the backbone models used in time series forecasting. It has been observed that linear-based models, such as Dlinear (Zeng et al., 2023), often exhibit superior performance compared to transformer-based models like Crossformer (Zhang and Yan, 2023), particularly in terms of computational efficiency and speed. These findings highlight the potential of structural optimizations in enhancing the capabilities of time series forecasting models. Nevertheless, the quest for an ideal model that balances performance, prediction accuracy, stability, and interpretability continues. The aim is to develop a model that can provide robust support for high-stakes scenarios where timely and reliable predictions are critical, thereby bridging the gap between theoretical advancements and practical application needs.
Evidence theory, also known as Dempster-Shafer theory, provides a mathematical framework for representing and combining degrees of belief in the presence of uncertainty (Shafer, 1976; DEMPSTER, 1967). It is particularly useful in situations where information is incomplete, ambiguous, or comes from diverse sources. The theory is grounded in the concept of a belief function, which quantifies the degree to which a hypothesis is supported by the available evidence, without necessarily committing to a probability distribution over the possible outcomes. One of the key strengths of evidence theory is its ability to handle the fusion of multiple information sources effectively. This makes it a powerful tool in fields such as data fusion, sensor networks, and multi-criteria decision-making, where diverse pieces of information need to be integrated to form a coherent union. In the realm of symbolic logic, evidence theory often appears as a reasoning framework that offers high interpretability. Unlike traditional probability theory, which requires precise probability assignments, evidence theory allows for a more nuanced representation of uncertainty. It does this through the use of belief and plausibility measures, which provide a range of values for the likelihood of an event occurring, rather than a single probability. An innovative application of evidence theory in machine learning is the development of evidence neural networks by Denoeux (Denoeux, 2000). These networks are designed for classification tasks and integrate the principles of evidence theory into the architecture of neural networks. By doing so, they enhance the network’s ability to handle uncertainty and provide a more robust and interpretable decision-making process. The evidence neural networks output belief masses that can be directly interpreted as measures of evidence supporting each class, thus offering a clear understanding of the network’s reasoning process.
Therefore, we introduce the Time Evidence Fusion Network (TEFN), a novel architecture designed specifically for regression tasks within the realm of time series forecasting. The conceptual underpinning of TEFN is the innovative strategy of dissecting the time series data into dual perspectives—time and channels—thereby treating each channel and temporal instance as a distinct and autonomous source of information. This approach allows us to leverage the principles of evidence theory to fuse these diverse information streams within an integrated framework, aiming to enhance the accuracy and reliability of predictions as depicted in Figure 1. In essence, TEFN is built upon the premise of simplicity and naivety in its design philosophy. It refrains from incorporating complex technical modifications that often render models opaque and difficult to interpret. Instead, TEFN maintains a straightforward structure that prioritizes clarity and ease of understanding, making it a transparent and accessible backbone for time series analysis. By adopting this minimalist approach, TEFN emerges as a refreshing alternative in the field, offering a new foundation upon which more sophisticated forecasting models can potentially be built, while retaining the interpretability and trustworthiness that are hallmarks of evidence theory.
In the architecture of TEFN, we introduce a pivotal component termed the Basic Probability Assignment (BPA) Module, which stands as the core of our evidence-theoretic approach to time series forecasting. The BPA Module operates on the fundamental principles of evidence theory, meticulously mapping various sources of information into mass distributions that are intrinsically linked to the desired target outcome. This mapping is achieved through the application of fuzzy membership functions, which assign membership grades to different fuzzy sets, thereby constructing a comprehensive target distribution as proposed by Deng (Deng, 2015). In contrast to the convolutional operations prevalent in image processing, which selectively condense features pertinent to the target through filters (Bengio et al., 1993), the BPA Module embodies an expansionary process, as illustrated in Figure 2. Rather than narrowing down to specific features, BPA embraces an expansive perspective, contemplating all potentialities within the event space, which is a collection of subsets derived from the sample space. This expansion across various events is particularly advantageous for the simplistic structures inherent in time series data, enabling a thorough exploration of the myriad hidden informational layers within. For the diverse distributions yielded by the multitude of information sources, we treat each source as an individual sample, and by amalgamating these samples, we utilize the expected forecasting values as the ultimate prediction, thereby delivering a robust and informed forecast that encapsulates the collective wisdom of all contributing sources.
@@
Figure 2. Comparison between convolution and basic probability assignment: convolution condenses information while basic probability assignment diverges to consider different possibilities.
In this paper, we delve into the practical nuances of the Time Evidence Fusion Network (TEFN), examining its precision, memory consumption, training velocity, sensitivity to parameters, and interpretability. To maintain equitable comparisons, all experimental data and configuration files are sourced from open-source initiatives. TEFN has partially achieved state-of-the-art (SOTA) performance on various real-world datasets, with a notable reduction in training duration and model complexity when contrasted with Transformer-based models such as Crossformer (Zhang and Yan, 2023) and Patch-based models like PatchTST (Nie et al., 2023). In terms of random hyperparameter selection, TEFN demonstrates minimal variability, affirming its stability. The contributions of this manuscript are summarized as follows:
  • The proposed method is well-suited for large datasets, showcasing exceptional performance.
@@
  • A novel evidential backbone BPA designed for extracting potential distributions from simple data structures is introduced.
@@
  • A BPA-based neural network, TEFN, is introduced for time series forecasting, outpacing prior research in terms of accuracy, efficiency, stability, and interpretability.
@@
  • The manuscript elucidates the incorporation of evidence theory concepts, such as the mass function, into the architecture of neural networks, offering a novel approach to uncertainty reasoning in machine learning.
This paper is organized as follows. We begin by reviewing relevant literature in the Section 2. Next, in the Section 3, we delve into the architecture and underlying mathematical principles of TEFN. In the Section 4, we present a series of experiments that demonstrate TEFN’s superior performance. Finally, we conclude the paper by summarizing our key findings in the Section 5.
2. Related Work
2.1. Dempster-Shafer Theory
@@
Dempster-Shafer theory, also known as Evidence Theory, operates on a different framework compared to traditional probability theory (Shafer, 1976; DEMPSTER, 1967). It relies on the concept of mass functions, which provide a more flexible representation of uncertainty, accommodating weaker constraints than traditional probabilities. When combining two mass functions, denoted as $m_{1}$ and $m_{2}$, the resulting fused mass function $m$ is determined using the Dempster-Shafer Rule (DSR), as described in Equation 1. The symbols $A$, $B$, and $C$ represent focal elements, which are typically combinations of all independent samples (or target labels, such as the classification targets).
(1) $m(A)=\sum_{B\cap C=A}m_{1}(B)*m_{2}(C)$
2.2. Basic Probability Assignment
BPA is a core concept in evidence theory, used to map real data $x$ into mass distributions. Mass distributions describe the degree of support for various hypotheses in the event space, providing a basis for reasoning and decision-making. The generation of BPA $m$ is often based on fuzzy logic, where fuzzy membership functions $\mu$ are used to map data points to different fuzzy sets, thereby constructing the target distribution in Equation 2 where $w$ and $b$ is weight and bias of the membership function. BPA can also be generated by assigning support degrees through Gaussian distributions $N$ (probability density function $f$) where $\mu$ and $\sigma$ are expectation and variance in Equation 3.
(2) $\begin{split}m=\mu(x)=wx+b\end{split}$
(3) $\begin{split}&x\sim N(\mu,\sigma)\\ &m=f(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}}\\ \end{split}$
The advantage of BPA lies in its ability to effectively handle uncertainty and improve the accuracy and reliability of predictions by fusing multi-source information.
2.3. Evidence Decision Making
@@
The process of Evidence Decision Making can be delineated into several sequential steps: basic probability assignment (BPA), evidence fusion, and decision making (DM) (Deng, 2015; Zhan et al., 2024). BPA serves as a pivotal initial phase, constituting a mapping from raw data to mass distribution, akin to the path from raw data to probability distribution. Subsequently, evidence fusion integrates distinct mass distributions from various data channels, culminating in a rationalized mass distribution. The most classic fusion method is the DSR (Shafer, 1976; DEMPSTER, 1967) mentioned in Section 2.1. Finally, decision making transpires as the progression from the attained distribution towards the designated target. The most classic decision-making method is to transform the mass distribution into the probability distribution of the target, called pignostic probability transformation (PPT) (Dubois et al., 1993) in Equation 4, where $x$ is a target label and $e$ are combinations of all possible target labels.
(4) $p(x)=\sum_{x\in e}\frac{m(e)}{(1-m(\emptyset)|e|)}$
Figure 3. The overall structure of TEFN
2.4. Linear Time Series Model
There are many linear time series forecasting models (LTSFM). LTSFM offer several advantages for modeling and forecasting time series data. Firstly, their simplicity and transparency make them easy to implement, interpret, and analyze. The parameters of linear models directly represent the relationships between input features and the output, providing insights into the underlying dynamics of the time series. Secondly, LTSFM exhibit high efficiency, requiring fewer computational resources and hyperparameters compared to complex models like transformer-based models (Vaswani, 2017; Zhang and Yan, 2023). This leads to faster training and inference times, making them suitable for real-time applications and large datasets.
A classic LTSFM model is Dlinear (Zeng et al., 2023). DLinear serves as a simple yet effective baseline for long-term time series forecasting. It is a model that employs a seasonal-trend decomposition technique to separate the trend and seasonal components of the time series data. Each component is then independently modeled using a single-layer linear regression. The final prediction is obtained by summing the outputs of the two linear regressions.
TEFN shares similarities with Dlinear in that both models utilize linear regression for prediction. However, there are key differences:
  • DLinear employs a simpler decomposition method compared to TEFN, which may limit its ability to capture intricate temporal dynamics.
  • DLinear focuses on capturing linear relationships and trends in the data, while TEFN focuses on the mining of internal hidden relationships, so it strengthens predictions through dimensionality increase and uncertain reasoning.
  • TEFN is a general framework that does not require identification and prediction of trends and periodicity, while Dlinear is more like a feature modeling prediction of time series.
3. Time Evidence Fusion Network
The structure of TEFN is inherently straightforward, achieving remarkable performance with a design featuring only two BPA modules, akin to employing just two convolutional kernels. We posit that the novel architecture of TEFN embodies the essence of a pure backbone model, stripped of domain-specific intricacies. This minimalist approach allows TEFN to focus solely on capturing the fundamental temporal patterns within the data, making it a versatile and adaptable foundation for various time series analysis tasks. The overall structure of TEFN is depicted in Figure 3.
@@ 3. Time Evidence Fusion Network
Figure 4. The operation process of a BPA module for a time series $T$ of length $L=5$. The sample space size in the BPA module is set as $|S|=2$, and the event space size is set as $|2^{S}|=2^{|S|}=4$. The operation of the BPA module is similar to that of the convolution module, but the BPA module diffuses different masses instead of summing them in the convolution.
3.1. Time Normalization and De-normalization
To expedite convergence towards local optima and enhance training efficiency, TEFN incorporates a normalization technique inspired by the Stationary model (Liu et al., 2022). This process computes the mean $\mu$ as shown in Equation 5 and variance $\sigma^{2}$ as shown in Equation 6 for each segment of the time series $x$. The subsequent normalization and de-normalization operations, presented in Equation 7 and Equation 8, respectively, act as symmetrical transformations that standardize the time series data. The neural network operation $Net(\cdot)$ is then applied to the normalized time series $x_{Norm}$, ensuring that the model learns from data with consistent scale and distribution. This normalization strategy effectively mitigates the influence of outliers and initial data scale, fostering faster convergence and improved stability during training.
(5) $\mu=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}$
(6) $\sigma^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}(x_{i}-\mu)^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}^{2}-\mu^{2}$
(7) $x_{Norm}=\frac{x-\mu}{\sigma}$
(8) $x_{De-Norm}=\sigma*\hat{y}_{Norm}+\mu=\sigma*Net(x_{Norm})+\mu$
3.2. Time Dimension Projection
In TEFN, we employ a linear layer $project(\cdot)$ to project dimensionality to an expansion dimensionality, transforming the input time series $x$ with a length of $L_{in}$ into a sequence $x^{\prime}$ with a length of $L_{in}+L_{pred}$ in Equation 9. Here, $L_{pred}$ represents the desired length of the prediction horizon. This operation is analogous to the time projection mechanism utilized in other neural network architectures, such as TimesNet (Wu et al., 2023).
(9) $x^{\prime}=project(x)=x\times W_{p}+b$
By extending the input vector to match the prediction length, the time projection layer facilitates the direct generation of future values based on the learned temporal patterns. This approach simplifies the model design and enables efficient prediction without the need for iterative or recursive computations, as commonly observed in autoregressive models.
3.3. Basic Probability Assignment Module
In TEFN, we leverage the principles of Evidence Theory to model the uncertainty and ambiguity inherent in time series data. According to Evidence Theory, each dimension of the time series can be represented as a mass function within a common event space. This event space is typically defined as the power set $2^{S}$ of the sample space $S$, where $S$ represents the possible values or labels for the data. For classification problems, $S$ corresponds to all the possible labels, while for regression problems, $S$ is treated as a set of anonymous samples, ensuring mutual exclusivity among them.
This representation allows us to capture the fuzzy nature of the data, where each dimension generates a mass distribution within the event space. To achieve this, we utilize parameterized fuzzy membership functions, as defined in Equation 10. Here, $D$ represents the dimensionality of the time series, encompassing both the time dimension $T$ and the channel dimension $C$. The indices $i$ and $j$ correspond to the specific dimensions within $D$, while $k$ represents the index of the event space dimension, denoted as $F$. The fuzzy membership function $\mu_{k}$ is defined in Equation 10 where $w_{D,j,k}$ and $b_{D,j,k}$ are the slope and intercept parameters of the membership function, respectively. These parameters are learned during training and allow the model to capture the varying degrees of membership for each data point within the event space.
(10) $m_{D,i,j,k}=\mu(x_{Norm,i,j})=w_{D,j,k}*x_{Norm,i,j}+b_{D,j,k}$
By employing evidence theory and fuzzy membership functions, TEFN effectively models the uncertainty in time series data and enables the generation of robust and accurate predictions. This approach provides a principled way to handle the inherent fuzziness and ambiguity in real-world time series, leading to improved performance and reliability in various applications.
To introduce non-linear characteristics into TEFN and enhance its modeling capabilities, we can modify the fuzzy membership function used in the BPA module. Instead of the linear form presented in Equation 10, we can adopt non-linear functions that capture more complex relationships within the data. Several options exist for non-linear membership functions, including:
  • Gaussian Function: This function models the data distribution as a bell-shaped curve such as Equation 3, capturing the likelihood of each data point belonging to different membership levels. The parameters of the Gaussian function can be learned to adapt to the specific characteristics of the time series data.
  • Sigmoid Function: This function smoothly transitions between 0 and 1, providing a flexible way to represent the membership of data points to different levels in Equation 11. The sigmoid function can capture both gradual and abrupt changes in membership values.
    (11) $\sigma(x)=\begin{cases}0&\text{if }x<0\\ \frac{1}{1+e^{-x}}&\text{if }x\geq 0\end{cases}$
  • Piecewise Functions: These functions consist of multiple linear or non-linear segments, allowing for more precise control over the shape of the membership function in Equation 12 where $g(x)$ is a non-linear function. This approach is useful for capturing complex patterns and relationships within the data.
    (12) $f(x)=\begin{cases}a_{1}x+b_{1}&\text{if }c_{1}x+d_{1}\leq 0\\ g_{1}(x)&\text{if }c_{2}x+d_{2}>0\text{ and }c_{3}x+d_{3}\leq 0\\ a_{2}x+b_{2}&\text{if }c_{4}x+d_{4}>0\\ \vdots&\\ a_{n}x+b_{n}&\text{otherwise}\end{cases}$
By utilizing non-linear membership functions, TEFN can effectively model the non-linear dynamics and interactions within the time series data. This enhancement enables TEFN to capture more subtle patterns and relationships, leading to improved accuracy and generalization in various time series forecasting tasks.
3.4. Expectation Fusion
TEFN addresses the challenge of handling multivariate time series data with two distinct dimensions: the time dimension $T$ and the channel dimension $C$. This results in the paralleled generation of different mass distributions for each dimension, denoted as $m_{C}$ and $m_{T}$. To integrate these mass distributions and produce a unified prediction, TEFN employs a fusion method based on expectations, as defined in Equation 13. Here, $y_{j,k}$ represents the fusion parameters, and $m^{\prime}_{i,j,k}$ corresponds to the transformed mass function for each dimension.
(13) $\hat{y}_{Norm,D,i,j}=E_{D,F}(y_{j,k})=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}y_{j,k}*m^{\prime}_{i,j,k}=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}m_{i,j,k}$
The fusion operation in Equation 13 can be interpreted as a linear transformation of the mass functions, where the summation of $y_{j,k}*m^{\prime}_{i,j,k}$ effectively combines the information from different dimensions. The expectation is then calculated by summing the transformed mass functions, resulting in a single mass function representing the fused prediction.
TEFN deliberately avoids using the Dempster-Shafer Rule of Combination (DSR) for several reasons. Firstly, the computational complexity of DSR is significantly higher compared to the expectation-based fusion method. Additionally, DSR involves point-to-point multiplication of mass distributions within the same dimension, which can lead to accuracy loss when dealing with high-dimensional data due to finite precision limitations.
Secondly, DSR is sensitive to extreme distributions, particularly those with a single element having a value of 1. This sensitivity can result in the final fused result being dominated by that single distribution, potentially leading to conflicts and loss of information from other distributions.
By employing the expectation-based fusion method, TEFN achieves a balance between computational efficiency and accuracy. This approach effectively integrates information from different dimensions while mitigating the risks of conflicts and accuracy loss associated with DSR. The final normalized prediction results are then aggregated across dimensions using the equation in Equation 14 and restored to their original scale through the de-normalization layer.
(14) $\hat{y}_{norm}=\sum_{D\in\{T,C\}}\hat{y}_{Norm,D}=\hat{y}_{Norm,T}+\hat{y}_{Norm,C}$
4. Experiments
The experimental section aims to comprehensively evaluate TEFN’s performance in long-term time series forecasting. Utilizing diverse datasets, TEFN will be benchmarked against established models, measuring accuracy, efficiency, and robustness. We will analyze its performance across different prediction horizons, assess its sensitivity to hyperparameters, and conduct ablation studies to understand the impact of individual components. Lastly, we will explore TEFN’s interpretability by analyzing the basic probability assignments generated by its BPA module.
4.1. Dataset
Our research employs the TEFN for the challenging task of long-term time series prediction. To ensure the model’s effectiveness and robustness across diverse scenarios, we meticulously selected five representative and extensively used datasets. These datasets encompass a wide range of real-world applications, providing a comprehensive evaluation of TEFN’s performance. The selected datasets are:
  • Electricity (ecl, [n.d.]): This dataset contains electricity load diagrams for a power grid from 2011 to 2014, offering insights into energy consumption patterns over time. It presents a challenging forecasting task due to its seasonal variations and potential anomalies.
  • ETT (4 subsets) (Li et al., 2021): This dataset comprises four subsets with varying lengths and complexities, representing diverse time series forecasting challenges. It allows us to assess TEFN’s ability to handle different data characteristics and prediction horizons.
  • Exchange (Wu et al., 2021): This dataset captures exchange rate fluctuations between major currencies. It presents a dynamic and volatile environment, testing TEFN’s ability to capture complex market trends and predict future exchange rates accurately.
  • Traffic (tra, [n.d.]): This dataset contains traffic flow data collected from sensors in a city. It reflects the temporal and spatial patterns of traffic congestion, providing a challenging task for forecasting traffic flow and congestion levels.
  • Weather (wea, [n.d.]): This dataset includes historical weather data, encompassing various meteorological parameters. It allows us to evaluate TEFN’s performance in predicting weather patterns, which are crucial for numerous applications like agriculture, disaster management, and transportation planning.
Each of these datasets is characterized by its large size and diverse features, offering a rich and challenging environment for evaluating TEFN’s capabilities. The extensive nature of the data ensures that the model is tested on a wide range of scenarios, providing valuable insights into its performance and generalizability. More dataset features are in the Table 1.
Table 1. The quantitative features of the dataset
Dataset Dimension Dataset Size
Electricity 321 (18317, 2633, 5261)
ETTm1, ETTm2 7 (34465, 11521, 11521)
ETTh1, ETTh2 7 (8545, 2881, 2881)
Traffic 862 (12185, 1757, 3509)
Weather 21 (36792, 5271, 10540)
Exchange 8 (5120, 665, 1422)
Table 2. This table presents a comprehensive comparison of prediction errors across various models for long-term time series forecasting tasks. The models are evaluated on four different prediction lengths: $L_{pred}\in\{96,192,336,720\}$, providing a thorough assessment of their performance under varying forecasting horizons. The color coding in the table highlights the ranking of each model’s performance across all datasets and prediction lengths, allowing for easy identification of the top-performing models.